# Advancing Generalizable Remote Physiological Measurement through the Integration of Explicit and Implicit Prior Knowledge

### v1

@@ IV Experiments
TABLE III: HR estimation results on MSDG protocol. $*{}^{*}$ means that these methods use the STMap as the input of CNN; ${}^{+}$ means that these methods are based on baseline (Rhythmnet [37] without GRU). The best results are shown in bold.
@@
$*{}^{*}$ [37] 5.79 7.91 0.78 7.39 10.49 0.77 3.38 5.17 0.84 8.97 12.16 0.49 10.16 14.57 0.34
$*{}^{*}$ [28] 5.55 7.62 0.79 7.24 10.27 0.78 3.41 5.23 0.84 8.88 11.69 0.50 10.04 14.44 0.35
$*{}^{*}$ [23] 5.43 7.71 0.80 7.23 10.25 0.78 3.69 5.48 0.81 8.45 11.64 0.51 10.01 14.35 0.36
${}^{*+}$ [44] 5.92 8.08 0.76 7.42 10.61 0.73 3.49 5.49 0.82 8.41 11.71 0.53 10.47 14.64 0.32
${}^{*+}$ [45] 5.73 7.97 0.78 7.69 10.83 0.78 3.41 5.21 0.83 8.35 11.67 0.54 9.94 14.29 0.36
${}^{*+}$ [57] 5.89 8.04 0.76 7.59 10.87 0.72 3.64 5.74 0.80 8.68 11.91 0.53 10.32 14.42 0.32
${}^{*+}$ [58] 5.59 7.68 0.81 7.24 10.14 0.78 3.27 5.01 0.86 8.37 11.62 0.54 9.82 14.16 0.37
${}^{*+}$ [59] 5.31 7.56 0.82 7.32 10.35 0.77 3.12 5.16 0.85 8.47 11.81 0.52 10.14 14.46 0.34
${}^{*+}$ [34] 4.67 6.79 0.86 6.71 9.59 0.81 2.88 4.69 0.89 7.86 11.15 0.58 9.27 13.79 0.41
$*{}^{*}$ 5.53 7.89 0.84 6.57 9.86 0.89 2.14 3.03 0.96 8.40 11.33 0.53 9.13 11.10 0.45
${}^{*+}$ w Ex-prior 4.48 6.56 0.88 5.46 8.58 0.88 1.93 2.60 0.97 7.39 10.78 0.63 8.90 10.73 0.46
${}^{*+}$ w Im-prior 4.72 6.89 0.85 5.64 9.05 0.87 1.78 2.48 0.97 8.07 11.33 0.54 8.81 11.31 0.50
${}^{*+}$ (Ours) 4.08 6.17 0.88 4.57 7.71 0.90 1.69 2.21 0.98 7.10 10.31 0.65 8.46 10.56 0.53
@@ IV-G Self-supervised learning
MoCo [63]
9.27 13.05 0.04
SIMSIAM [64]
8.43 11.73 0.14
BOYL [65]
8.98 12.43 0.08
SIMCLR [66]
8.57 11.94 0.10
Gideon21 [40]
9.80 15.48 0.38
Contrast-phys [27]
8.55 12.65 0.40
rPPG-MAE [7]
7.83 11.19 0.48
@@ IV-H Semi-supervised learning
HTER_test normal-$\downarrow$ EER normal-$\downarrow$ AUC normal-$\uparrow$ HTER_test normal-$\downarrow$ EER normal-$\downarrow$ AUC normal-$\uparrow$

### v2

@@ IV Experiments
TABLE III: HR estimation results on MSDG protocol. means that these methods use the STMap as the input of CNN; + means that these methods are based on baseline (Rhythmnet [37] without GRU). The best results are shown in bold.
@@
Rhythmnet [37] 5.79 7.91 0.78 7.39 10.49 0.77 3.38 5.17 0.84 8.97 12.16 0.49 10.16 14.57 0.34
Dual-GAN [28] 5.55 7.62 0.79 7.24 10.27 0.78 3.41 5.23 0.84 8.88 11.69 0.50 10.04 14.44 0.35
BVPNet [23] 5.43 7.71 0.80 7.23 10.25 0.78 3.69 5.48 0.81 8.45 11.64 0.51 10.01 14.35 0.36
AD∗+ [44] 5.92 8.08 0.76 7.42 10.61 0.73 3.49 5.49 0.82 8.41 11.71 0.53 10.47 14.64 0.32
GroupDRO∗+ [45] 5.73 7.97 0.78 7.69 10.83 0.78 3.41 5.21 0.83 8.35 11.67 0.54 9.94 14.29 0.36
Coral∗+ [57] 5.89 8.04 0.76 7.59 10.87 0.72 3.64 5.74 0.80 8.68 11.91 0.53 10.32 14.42 0.32
VREx∗+ [58] 5.59 7.68 0.81 7.24 10.14 0.78 3.27 5.01 0.86 8.37 11.62 0.54 9.82 14.16 0.37
NCDG∗+ [59] 5.31 7.56 0.82 7.32 10.35 0.77 3.12 5.16 0.85 8.47 11.81 0.52 10.14 14.46 0.34
NEST∗+ [34] 4.67 6.79 0.86 6.71 9.59 0.81 2.88 4.69 0.89 7.86 11.15 0.58 9.27 13.79 0.41
Baseline 5.53 7.89 0.84 6.57 9.86 0.89 2.14 3.03 0.96 8.40 11.33 0.53 9.13 11.10 0.45
Greip∗+ w Ex-prior 4.48 6.56 0.88 5.46 8.58 0.88 1.93 2.60 0.97 7.39 10.78 0.63 8.90 10.73 0.46
Greip∗+ w Im-prior 4.72 6.89 0.85 5.64 9.05 0.87 1.78 2.48 0.97 8.07 11.33 0.54 8.81 11.31 0.50
Greip∗+ (Ours) 4.08 6.17 0.88 4.57 7.71 0.90 1.69 2.21 0.98 7.10 10.31 0.65 8.46 10.56 0.53
@@ IV-G Self-supervised learning
MoCo [63] 9.27 13.05 0.04
SIMSIAM [64] 8.43 11.73 0.14
BOYL [65] 8.98 12.43 0.08
SIMCLR [66] 8.57 11.94 0.10
Gideon21 [40] 9.80 15.48 0.38
Contrast-phys [27] 8.55 12.65 0.40
rPPG-MAE [7] 7.83 11.19 0.48
@@ IV-H Semi-supervised learning
HTER_test $\downarrow$ EER $\downarrow$ AUC $\uparrow$ HTER_test $\downarrow$ EER $\downarrow$ AUC $\uparrow$